Displaying reports 261-280 of 1079.Go to page Start 10 11 12 13 14 15 16 17 18 End
Reports until 14:36, Friday 13 December 2024
Equipment Loans
Torrey Cullen - posted 14:36, Friday 13 December 2024 - last comment - 09:02, Monday 16 December 2024(12045)
350-2000nm IR viewer loan from cryo lab

[Torrey, JC]

JC and I picked the lock of the locked lista cabinet in the cryo lab and got the IR viewer. I used it to observe the power distribution center. I did not observe any significant amounts of scattered 1550nm light. There was a fair amount visible on the SHG sled where irises are being used to block the 0th order beam but this is 775. The amount of power there is known to be low though, I don't think this is an issue. 

In short, I'm not sure we need one of these viewers. I could not observe any scatter where I expected there to be some. It could easily see 775 nm, as expected as they are several orders of magnitude more sensative at this wavelength. One caveat potentially is this looks like a much older model than what is currently being sold, maybe the technology has improved since this older version.

Images attached to this report
Comments related to this report
Ian MacMillan - 09:02, Monday 16 December 2024 (12049)

This loan has been posted to the Adhikari elog. Please update when it is returned.

GQuEST General
Daniel Grass - posted 12:33, Thursday 12 December 2024 (12044)
Cart for SNSPD Dewer Assembled with Custom Plates

[Alex, Sander, Daniel]
 

The Thorlabs parts to make a cart with a breadboard top did not fit together. In addition, the POC001 cart could shake pretty easily with the included wheels. I therefore designed some custom plates, one set to attach new wheels from McMaster and another set to interface from the PWA075 to the B2436G Breadboard. To attached the plate to the cart for the wheels, I used Stainless Steel 15 mm long, button head M6 screws. To attached the McMaster Wheels to the plate, I used 3/4" long, 5/16-24 screws with 5/8" OD washers.

To attach the plate to the breadboard, I used a 3/4" long 1/4-20 button head screw with a lock washer (similar to this product) to prevent the plate from rotating (parts held by a single screw without a lock washer tend to rotate). To hold the PWA075 to the breadboard, I used some Thorlabs stainless steel 5/8" long, 1/4-20 screws with some washers.

To attach the PWA075 to the cart, I used a 55 mm long M12 set screw that I screwed into the PWA075 since that was a blind hole. We used 8 rotations to put the set screw into the cart. I used nuts and washers to prevent PWA075 and the cart from spinning on the set screw.

When possible, I used alloy steel fasteners for its strength. The parts from McMaster were fairly greasy, so I washed them in simple green in DI water (1:30 ratio), sonicated for 10 minutes, and washed in DI water.

I made the plates out of Aluminum 6061. Attached are the part files, STEP files, ProtoMax Waterjet machining files, and drawings for making the parts in the mill. I had to machine 2 plates that interface with the breadboard as drawn and 2 parts with the opposite face milled because the parts need the be flipped accross the axes transverse to the breadboard face. The plates were greasy, so I washed them in simple green in DI water (1:30 ratio), sonicated for 3 minutes, and washed in DI water.

We assembled everything together, and everything fits. See attached photo.

Images attached to this report
Non-image files attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 16:02, Wednesday 11 December 2024 (12043)
Moku Meeting

Jeff and I met with the Moku CEO and several employees today. Meeting went well. One take away I wanted to make public knowledge is the different capabilities of the computer app vs the ipad app. Attached are some example screenshots. moku1.png can take a spectrum and a transfer function of the same signal at the same time, something I have not been doing up until this point. moku2.png as a starter point for when we are chaining cavities together in the future. 

Images attached to this report
Environmental General
Ian MacMillan - posted 14:05, Tuesday 10 December 2024 (12042)
Very High Lab Humidity

[Ian, Torrey]

We noticed that it is very humid in the lab today. It is also much warmer than usual. The attached plots for the RbQ lab (A) and the GQuEST lab (B) show a spike in humidity between the 8th and 9th of December from about 20% relitive humidity to about 50% relitive humidity. The water does not seem to be condensing.

Non-image files attached to this report
Computing General
Jeffrey Wack - posted 11:39, Tuesday 10 December 2024 (12041)
advligorts setup on babbage

I have followed this guide to set up software on the CyMAC front-end machine running bullseye (Debian 11), also known as babbage.

Notes/Questions:

Kernel Module Status
mbuf               ***NOT LOADED**
gpstime            ***NOT LOADED**
rts_cpu_isolator   ***NOT LOADED**
rts_logger         ***NOT LOADED**
Kernel Module Status = SOME MISSING


System Status
system         epics          module         awgtpman
------------------------------------------------------------
Systems Status =       ALL ACTIVE


Streamer Status
rts-local_dc       OFF
rts-transport@cps_xmit OFF
Streamer Status =      DEGRADED

Overall Status =       DEGRADED

Vacuum Systems General (GQuEST)
Daniel Grass - posted 17:08, Monday 09 December 2024 (12040)
3 Laser Filter Cavity Vacuum Cube Flanges Moved and Added

[Jeff, Daniel]

I moved the "output" Laser Filter Cavity Cube closer to the edge of the table so that we could work on it more easily. Jeff and I took off the 10" CF flange with an angled 2.75" CF flange off the cube, wrapped it in foil, and put it away with the other holometer parts. I didn't like this flange as it stuck out past the custom aluminum base and didn't seem very useful. In its place, we moved the 10" CF flange with two 15 pin D-Subs from the top to this location (this flange will be facing South when the cube is placed back). We moved a 10" to 4.5" zero length reducing flange from the side to the top. In that place, we added a new 10" to 6" zero length reducer. I believe we match up the hole pattern so that the 6" to 2x2.75" reducer will have the windows 5" above the cube bottom as designed.

I cleaned parts of the cube and the flanges with the Wipes rated for cleaning UHV, but the surfaces were pretty clean to begin with.

We only tightened the screws so that the washers couldn't move. I still need to tighten the screws to compress the copper gasket.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 18:24, Thursday 05 December 2024 (12038)
Holometer Parts Disassembled for Use in the Laser Filter Cavity

[Jeff, Daniel]
 

We dissassembled the 8" to 6" reducing tee from a ~20" long tube (with 6" CF Flange ends) and wrapped the open ends in UHV foil. We also took off two 6" CF flange gate valves from two of these tubes and wrapped the open ends in foil. We put everything except the reducing tee away into the white cabinet with other Holometer equipment. The reducing tee is on the table by the central vessel.

Controls General
Jeffrey Wack - posted 17:28, Thursday 05 December 2024 (12037)
Initial piezo lock of output filter cavity #2

[Jeff, Torrey]

We performed an initial lock of the second output filter cavity, and we eliminated the the 890Hz noise issue by putting the LNA10 Oscilloscope Preamplifier as an input buffer to the seeder DC modulation input.

Laser lock

The qualily of the lock with the laser is approximately as good as in B102. There is a 115kHz resonance with a high gain controller.

Piezo lock - no amplifier

We performed a lock with no amplifier by manually scanning the laser until we saw flashes, then activating the piezo loop. This loop was limited by a 2.6kHz resonance with a high gain controller

Piezo lock - with amplifier

The EVAL-ADHV4702-1 provides a gain of 21V/V, allowing us to scan on the piezo and lock. The 15V power supplies from Amazon are very bad and result in unacceptable 120Hz noise in the amplifier output. We switched to using a Newport 15V supply and this problem was eliminated. The lock with the amplifier is of a comparable quality to the laser lock, but it is limited by a 25kHz resonance. We will take data soon to better characterize the loop.

 

Computing General
Jeffrey Wack - posted 12:03, Thursday 05 December 2024 (12036)
SSHing into front-end machines

Both cymac front-end machines have been installed in the server cabinet in the control room (B111B east) with an ethernet connection to the wall and have openssh-server installed. They have been added to the local DNS, so you can ssh into them using 

ssh cymac@turing.lab

ssh cymac@babbage.lab

I have also added the mini PC in the control room, "brewster" to the DNS, but cannot manage to ssh into it.

Equipment General
Torrey Cullen - posted 09:45, Thursday 05 December 2024 - last comment - 17:16, Friday 06 December 2024(12034)
Power Meter Repair

[Ian, Torrey]

While working in the lab, the ND filter attached to the power meter would not slide into place. Upon inspection we noticed that the sensor was slightly askew . We took it apart and found that the glue used to glue the sensor to the plastic back had come apart (they used a miniscule amount for some reason). We glued the sensor back to the plastic casing and cleaned the ND filter. Put things back together and things are working has intended.

Additionally we went to look at the second power meter Alex uses and the ND slider was not working on that one either. We will repair this one as well, have not gotten to it yet.

Images attached to this report
Comments related to this report
Torrey Cullen - 17:16, Friday 06 December 2024 (12039)

Took apart and re-glued the one alex uses. It is drying under some alluminum foil in B111B for the time being. Will reassemble Monday.

Computing General
Jeffrey Wack - posted 16:03, Wednesday 04 December 2024 (12032)
addendum to BIOS settings for front-end machines

[Jeff, Ian]

continuing from [11998]

After following this procedure to configure the BIOS settings, the computer did not boot. This was because during step #1, resetting all the settings, the boot order was modified so that 'Boot Option #1' was set to [Hard Disk: (Bus 01 Dev 00) PCI RAID Adapter]. After changing 'Boot Option #1' to [UEFI Hard Disk:debian], the computer boots as expected.

Vacuum Systems General (GQuEST)
Daniel Grass - posted 22:18, Tuesday 03 December 2024 (12031)
Power Recycling Cavity (PRC) and Demonstrator Interferometer (IFO) 10 in Vacuum Cubes Placed on Optics Table

[Torrey, Ian, Daniel]

We placed the PRC and Demonstrator IFO Vacuum Cubes on the optics table and secured them to the table with  1" long, 1/4-20 screws on each corner. To do this, we placed the cubes in the Northwest corner of the optics table and pushed them into place. For the PRC, we used some used clean room wipes to reduce the friction. The IFO was close enough that we just pushed it. The aluminum is able to slide on the breadboard somewhat easily. 

Torrey and I then removed the remaining 5" long, 8" diameter bellows and a 10" to 8" zero length reducer flange from the IFO vacuum cube so that the OFC 1 sled didn't have it in the way. We covered the exposed end of the 10" to 8" zero length reducer flange and the cube in foil.

It's now time to start replacing flanges. I added a tab on this google sheet to track the 10" flanges. It seems that we have almost everything we need to make a functional Laser Filter Cavity (LFC) and demonstrator IFO vacuum chambers. I think we need 2 more CF Flange 2.75" 1550 nm AR Coated Windows for the IFO.

Controls Optimal Controls
Ian MacMillan - posted 17:51, Tuesday 03 December 2024 (12030)
Plot Upgrades When Making a System in Buzz

I updated some of the plots in the buzz code for the ASC Solver folder when making a system as described in the previous log post. The new plots come with color coordination so that it is easy to see which lines should overlap. The first plot (SPOFF_diagnostic_bode.pdf) shows the large system plotted against the original parts that were used to make large system. The short story of this plot, and rally both of the plots, is that all of the same color lines should overlap. When each part is added to the larger system it should still have the same frequency response. This plot offers a quick look at all of the parts to make sure the decomposition of the parts is working correctly in the makeSPOFF() function. The second plot (FOM_bode_from_SPOFF.pdf) does a very similar thing but with the FOMs where it plots the individual state space systems of the FOMs then puts them into the large SPOFF and plots them again from there. Ideal these should line up which they do. The third green line on those plots shows the FOMs that have been pulled out of the SPOFF after the SPOFF has been balanced using the ssutil.balance_sys_gain() function. This is to show how the gain balancing is affecting the overall numerical limitations of the system. You can see in the BNS FOM that in the balanced system the BNS FOM can not go as low in magnitude. It hits a numerical wall earlier than the un balanced versions of the BNS FOM. 

I tried to run the HiB solver with the pre-balanced version. (i.e. balanced before it was given to the solver) but is showed similar results with one exception. The range calculated from the closed loop PSD seemed to give less wild values. As you can see in the third figure (rms_HiB_range_from_PSD.pdf). The values dont make sense but they at lease don't seem random.

Non-image files attached to this report
Equipment General
Alex Ramirez - posted 15:18, Tuesday 03 December 2024 (12029)
Thorlabs Cart for Dewar Update

We have recieved our thorlabs cart but encountered a few hiccups that Daniel and I will address:

  1. The cart seems well crafted, but the casters it is on allow the table to wobble and sway as they lock from moving but are free rotating. To address this, we are looking to retrofit 8020 leveling casters on the cart to allow for easy movement but firm ground contact when set in place. T
    1. These are the casters and attached is an image of them with a custom baseplate that Daniel will help me machine to mount to the cart. https://8020.net/2715.html
  2. Upon further investigation, the passive breadboard isolation mounts do not fit with this cart (although I had specifically asked the Thorlabs engineer/technician about them prior to purchasing). Daniel and I are on the case for making an adapter to allow for them to be used in place of the mounting posts it comes with which do not mechanically secure the optic table to the cart in any way, it just lays on top. So we will use the nice damping mounts and create the brace we need to get the tabletop securely mounted to the cart via the dampers. 

    Further work relating to the mounting of the damping pads will be added when finalized.
Images attached to this report
GQuEST Output Filter Cavities
Torrey Cullen - posted 11:17, Tuesday 03 December 2024 - last comment - 14:11, Tuesday 03 December 2024(12027)
Work in Lab notes 12/3

Notes from working in lab:

-With the amplifier not amplifying (just nominal ~300 mW output state) the REFL 1550 PD for OFC1 needs a lower ND filter on it. I've swapped it. Note that this should be swapped back BEFORE turning the amplifier on. I've put a note on the amplifier as a reminder.

-The error signal size is quite small. Had an idea from something Lee said a while ago. When we were setting up the new space someone saw the tank circuit box hanging off the EOM and suggested adding a small cable to it. Increasing the length between the inductor and the EOM significantly decreased the error signal amplitude. I've removed the cable. (pics attached)

-In the move I mixed up which tank circuit is for which EOM. They have been identified and labeled. Both tank circuits are on their respective EOMs.

-The error amplitude (with +24 db input boost on the laser lock box) is now ~100 mV and the RMS of the error signal when the cavity is laser locked is 10 mV.

-Went to start 775 OFC1 alignment but the AOM for this path is not deflecting a beam now. I don't know the cause.

Images attached to this report
Comments related to this report
Torrey Cullen - 14:11, Tuesday 03 December 2024 (12028)

Focusing on other things until AOMs are diagnosed. Afternoon work:

-PDs are powered for OFC2

-BNCs are plugged into patch panel and Moku 2.

-Looking for flashing in OFC2, can change which moku scans the laser freq by swapping patch panel locations. Much easier than before.

-775 path EOM has been optimized to give max error amplitude.

-OFC2 has been recovered and is fully functional (both paths).

Plan is to fix the AOM for OFC1, align 775 for OFC1, and then start thinking about OFC3. In B102 my moku systems were a little wonky, with the patch panel and recabling from scratch each filter cavity has its own moku now. OFC1 = Moku 1 / OFC2 = Moku 2.

GQuEST Output Filter Cavities
Torrey Cullen - posted 12:34, Monday 02 December 2024 - last comment - 16:59, Monday 02 December 2024(12024)
B111 Lab Progress

[Torrey, Ian]

General progress notes:

-The three photo diodes required for OFC1 activities are powered.

-All cables to and from the moku/patch panel are set up for OFC1 activities, including PDs, EOM, and DC modulation for the laser.

-1550 path is realigned and can be locked with the laser.

 

Problem:

The ~900 Hz signal that arrises from having a cable plugged into the DC modulation port of the laser needs to be solved. This is a known problem from when we were in B102 but it seems much worse over here. Here is what I know. 

1) Light is blocked. DC modulation port plugged in. ~890 Hz signal observed on PD. This is not optical and must be electrical.

2) Light blocked. DC modulation port is unplugged from the patch panel. Oscillations are gone. To clarify, I am unplugging it here: PXL_20241202_193738276~2.jpg

3) This interferes with the error signal enough to not give a good laser lock. 

4) The oscillations can be seen when the channel is DC coupled and light is hitting the PD.

5) Both the laser and the photo detector are plugged into a blue outlet, but not the same blue outlet.

If anyone has any insights on this let me know. I am curious to see if the PD is plugged into a white outlight if this makes a difference. Lee has suggested a type of circuit to fix this potential grounding issue in the past although I am blanking on the name of it. 

Images attached to this report
Comments related to this report
Torrey Cullen - 16:59, Monday 02 December 2024 (12026)

After some frustrating moku glitches and then some realignment, the second output filter cavity now has 775 light. Pending both 775 paths, and 1550 in OFC2, we should be ready for filter cavity science again.

Lab Infrastructure General
Daniel Grass - posted 19:38, Tuesday 26 November 2024 - last comment - 17:17, Monday 16 December 2024(12023)
USB Cable from Vacuum Oven in B110 to B111B Control Room Computer Set Up

[Sander, Daniel]

We routed a USB Extension Cable from the vacuum oven in B110 through B111D to the B111B control room. The cable trays above the mobile clean rooms in B111B were very hard to access. Maybe we should get some sort of cheap grabber tool from Amazon.

The cables make it to the computers in B111B but just barely. A ~6 ft USB extension cable would ensure there is plenty of slack in the system.

The computer recognized that something was plugged in, so it seems the extension cables work. Next week, I will finish setting up the vacuum oven software that I downloaded. I previously got this software to work on my laptop.

Comments related to this report
Daniel Grass - 16:19, Monday 02 December 2024 (12025)

The computer is able to communicate with the vacuum oven after very minimal software setup.

Daniel Grass - 17:17, Monday 16 December 2024 (12052)

I added another 50" USB extension so that there isn't tension on the cables. This is ~45" too long, but the extra cable is not in the way and the oven is able to communicate with the computer and vice versa.

Computing General
Jeffrey Wack - posted 12:23, Tuesday 19 November 2024 - last comment - 16:05, Wednesday 04 December 2024(11998)
BIOS settings for Turing (front-end machine)

I am following this DCC document to configure the BIOS settings of Turing, one of our front-end machines. 

Machine info

sudo dmidecode -s baseboard-product-name

X11SRL-F

sudo dmidecode -s processor-version

Intel (R) Xeon(R) W-2245 CPU @ 3.90GHz

The BIOS version is 2.8.

BIOS settings

First, I reset all BIOS settings to default using 'Restore Defaults' in the 'Save & Exit' menu.

Advanced > CPU Configuration: Hyper-Threading [ALL] - Disabled

Advanced > CPU Configuration > Advanced Power Management Configuration > CPU P State Control: SpeedStep (Pstates) - Disabled

Advanced > CPU Configuration > Advanced Power Management Configuration > CPU C State Control: Enhanced Halt State (C1E) - Disabled

Advanced > CPU Configuration > Advanced Power Management Configuration > Package C State Control: Package C State - C0/C1 state

Advanced > PCIe/PCI/PnP Configuration: Above 4G Decoding - Enabled

IPMI > BMC Network Configuration: Update IPMI LAN Configuration - Yes

IPMI > BMC Network Configuration: Configuration Address Source - Static

IPMI > BMC Network Configuration: IPv6 Support - Disabled

 

I will need to follow up about the Station IP address. What is the local CDS admin LAN? Will we be doing this? For now I have set the static IP address to be the same address that was DHCP assigned.

 

Comments related to this report
Jeffrey Wack - 16:05, Wednesday 04 December 2024 (12033)

NOTE: make sure that debian is selected as Boot Option #1 in the boot settings, this may be changed by resetting all settings

Displaying reports 261-280 of 1079.Go to page Start 10 11 12 13 14 15 16 17 18 End